# Differentially Private Bilevel Optimization

### v2

@@ 2.1 Setting
We impose the following assumptions, all of which are standard in the bilevel optimization literature.
@@
    $\mathcal{X}\subseteq\mathbb{R}^{d_{x}}$ is a closed convex set.
@@
    $F(\mathbf{x}_{0})-\inf_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\leq\Delta_{F}$ for some initial point $\mathbf{x}_{0}\in\mathcal{X}$.
@@
    $f$ is twice differentiable, and $L_{1}^{f}$-smooth.
@@
    For all $\xi\in\Xi:~{}f(\cdot,\cdot\,;\xi)$ is $L_{0}^{f}$-Lipschitz (hence, so is $f$).
@@
    $g$ is $L_{2}^{g}$-Hessian-smooth, and for all $\mathbf{x}\in\mathcal{X}:~{}g(\mathbf{x},\cdot)$ is $\mu_{g}$-strongly-convex.
@@
As mentioned, these assumptions are standard in the study of bilevel problems and are shared by nearly all of the previously discussed works. In particular, the strong convexity of $g(\mathbf{x},\cdot)$ ensures that $\mathbf{y}^{*}(\mathbf{x})$ is uniquely defined, which is generally required in establishing the regularity of the hyporobjective. Indeed, it is known that dropping this assumption, can, in general, lead to pathological behaviors not amenable for algorithmic guarantees (cf. Chen et al. 2024; Bolte et al. 2025 and discussions therein). For the purpose of differential privacy though, the strong convexity of $g(\mathbf{x},\cdot)$ raises a subtle issue. As the standard assumption in the DP optimization literature is that the component functions are Lipschitz, which allows privatizing gradients using sensitivity arguments, strongly-convex objectives cannot be Lipschitz over the entire Euclidean space. [^3] If $g(\mathbf{x},\cdot\,;\xi)$ were Lipschitz over $\mathbb{R}^{d_{y}}$ for all $\xi\in\Xi$, then so would $g(\mathbf{x},\cdot)$, contradicting strong convexity. Therefore, strongly-convex objectives are regularly analyzed in the DP setting under the additional assumption that the domain of interest is bounded. For bilevel problems, the domain of interest for $\mathbf{y}$ is the lower level solution set, thus we impose the following assumption.
@@
Following Assumptions 2.5 and 2.6, we denote $\ell:=\max\{L_{0}^{f},L_{1}^{f},L_{0}^{g},L_{1}^{g},L_{2}^{g}\},~{}\kappa:=\ell/\mu_{g}$.
@@ 3 Warm up: Privacy can leak from lower to upper level
In this section, we provide an illuminating example of privacy breaking in first-order bilevel optimization, which motivates the algorithmic approach presented later in the paper. In particular, we exemplify that even when the upper level problem supposedly does not depend on the dataset at all, the bilevel structure can give rise to hypergradients that are non-private.
@@
Given any dataset $S=\{\xi_{1},\dots,\xi_{n}\}\subset\mathbb{R}^{d}$, consider the bilevel problem corresponding to the objectives over $\mathbb{R}^{2d}:$
@@
These are smooth non-negative objectives, and $g$ is strongly-convex with respect to $\mathbf{y}$, so we see that Assumptions 2.5 and 2.6 hold. Further note that the upper-level objective $f$ does not depend on the data $S$ at all. Nonetheless, we will show that a single hypergradient computation can break privacy.
To see this, note that for any $\mathbf{x}\in\mathbb{R}^{d}:~{}\mathbf{y}^{*}(\mathbf{x})=\arg\min_{\mathbf{y}}g(\mathbf{x},\mathbf{y})=\frac{1}{n}\sum_{i=1}^{n}\xi_{i}$ since the squared loss is minimized by the average. Furthermore, $\nabla_{x}f(\mathbf{x},\mathbf{y})=\mathbf{x}+\mathbf{y}$ and $\nabla^{2}_{xy}g\equiv\mathbf{0}$, so by Eq. (1) we get that
$\nabla F(\mathbf{x})=\mathbf{x}+\mathbf{y}^{*}(\mathbf{x})=\mathbf{x}+\frac{1}{n}\sum_{i=1}^{n}\xi_{i}~{}.$
In particular, computing the hypergradient at the origin simply returns the dataset’s mean $\nabla F(\mathbf{0})=\frac{1}{n}\sum_{i=1}^{n}\xi_{i}$, which is well-known to be a non-private statistic of the dataset $S$ (cf. Kamath et al., 2019).
The algorithms to follow incorporate a fix to overcome such privacy leaks: instead of ever computing $\mathbf{y}^{*}(\mathbf{x})$ (as non-private algorithms do up to exponentially small precision), we estimate it using an auxiliary private method, and use the resulting private point $\widetilde{\mathbf{y}}^{*}(\mathbf{x})$ to estimate the hypergradient at $\mathbf{x}$. This leads to non-negligible bias in the hypergradient estimators compared to non-private first-order bilevel optimization, that we need to account for in our analysis. We further discuss this in Section 4.1.
@@ 4 Algorithm for DP bilevel ERM
In this section, we consider the ERM bilevel problem, namely (BO/ERM), for which we denote the hyperobjective by $F_{S}$. Our algorithm is presented in Algorithm 1. We prove the following result:
@@
Assume 2.5 and 2.6 hold, and that $\alpha\leq\ell\kappa^{3}\min\{\frac{1}{2\kappa},\frac{L_{0}^{g}}{L_{0}^{f}},\frac{L_{1}^{g}}{L_{1}^{f}},\frac{\Delta_{F}}{\ell\kappa}\}$. Then there is a parameter assignment $\lambda\asymp\ell\kappa^{3}\alpha^{-1},~{}\sigma^{2}\asymp\ell^{2}\kappa^{2}T\log(T/\delta)\epsilon^{-2}n^{-2},~{}\eta\asymp\ell^{-1}\kappa^{-3},~{}T\asymp\Delta_{F}\ell\kappa^{3}\alpha^{-2}$, such that running Algorithm 1 satisfies $(\epsilon,\delta)$-DP, and returns $\mathbf{x}_{\mathrm{out}}$ such that with probability at least $1-\gamma:$
@@
where $K_{1}=\mathcal{O}(\Delta_{F}^{1/4}\ell^{3/4}\kappa^{5/4}),~{}K_{2}=\mathcal{O}(\Delta_{F}^{1/6}\ell^{1/2}\kappa^{11/6})$.
@@
5:     $\mathbf{x}_{t+1}=\arg\min_{\mathbf{u}\in\mathcal{X}}\left\{\left\langle\widetilde{\mathbf{g}}_{t},\mathbf{u}\right\rangle+\frac{1}{2\eta}\left\|\mathbf{u}-\mathbf{x}_{t}\right\|^{2}\right\}$triangleright If $\mathcal{X}=\mathbb{R}^{d_{x}}$, then $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta\widetilde{\mathbf{g}}_{t}$
@@
7:$t_{\mathrm{out}}=\arg\min_{t\in\{0,\dots,T-1\}}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|$.
@@ 4.1 Analysis overview
Lemma 4.3 (Kwon et al. 2023; Chen et al. 2024).
If $\lambda\geq 2L_{1}^{f}/\mu_{g}$, then:
  1. (i)
    $\left\|\mathcal{L}^{*}_{\lambda}-F\right\|_{\infty}=\mathcal{O}(\ell\kappa/\lambda)$.
  2. (ii)
    $\left\|\nabla\mathcal{L}^{*}_{\lambda}-\nabla F\right\|_{\infty}=\mathcal{O}(\ell\kappa^{3}/\lambda)$.
  3. (iii)
    $\mathcal{L}_{\lambda}^{*}$ is $\mathcal{O}(\ell\kappa^{3})$-smooth (independently of $\lambda$).
@@
This strategy raises a privacy concern, as we exemplified in Section 3: Since $\mathbf{y}^{*}(\mathbf{x}),\mathbf{y}^{\lambda}(\mathbf{x})$ are required in order to compute $\nabla\mathcal{L}_{\lambda}^{*}(\mathbf{x})$, and are defined as the minimizers of $g(\mathbf{x},\cdot),\mathcal{L}_{\lambda}(\mathbf{x},\cdot)$ which are data-dependent, we cannot simply compute them under DP. In other words, even deciding where to invoke the gradient oracles, can already leak user information, hence breaking privacy before the gradients are even computed. We therefore must resort to approximating them using an auxiliary private method, for which we use DP-Loc(alized)-GD (Algorithm 2).444 We can replace the inner solver by any DP method that guarantees with high probability the optimal rate for strongly-convex objectives, as we further discuss in Appendix A.
In our analysis, we crucially rely on the fact that $g(\mathbf{x},\cdot)$ $\mathcal{L}_{\lambda}(\mathbf{x},\cdot)$ are both strongly-convex, which implies that optimizing them produces ${\widetilde{\mathbf{y}}}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t}$ such that the distances to the minimizers, namely $\left\|{\widetilde{\mathbf{y}}}_{t}-\mathbf{y}^{*}(\mathbf{x}_{t})\right\|,\left\|\widetilde{\mathbf{y}}^{\lambda}_{t}-\mathbf{y}^{\lambda}(\mathbf{x}_{t})\right\|$, are small. The distance bound is key, since then Eq. (3) allows using the smoothness of $f,g$ to translate the distance bounds into an inexact (i.e. biased) gradient oracle for $\nabla\mathcal{L}_{\lambda}^{*}(\mathbf{x}_{t})$, computed at the private points ${\widetilde{\mathbf{y}}}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t}$. We therefore are able to bound the gradient bias induced by privatizing the inner solvers, as follows:
@@
If $\lambda\geq\max\{\frac{2L_{1}^{g}}{\mu_{g}},\frac{L_{0}^{f}}{L_{0}^{g}},\frac{L^{f}_{1}}{L^{g}_{1}}\}$, then with probability at least $1-\gamma$, for all $t<T:$
@@
Bounding the Lipschitz constant of $\mathcal{L}_{\lambda,i}^{*}$ naively by applying the chain rule would result in the bound $\mathrm{Lip}(\mathbf{y}^{*})(L_{0}^{f}+2\lambda L_{0}^{g})\lesssim\lambda\mathrm{Lip}(\mathbf{y}^{*})L_{0}^{g}$, where $\mathrm{Lip}(\mathbf{y}^{*})$ is the Lipschitz constant of $\mathbf{y}^{*}(\mathbf{x}):\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{y}}$. Unfortunately, this bound grows with the penalty parameter $\lambda$, which will eventually be set large, and in particular, will grow with the dataset size $n$. We therefore need to derive a more nuanced analysis that allows obtaining a significantly tighter Lipschitz bound. We achieve this by showing that $\|\mathbf{y}^{\lambda}(\mathbf{x})-\mathbf{y}^{*}(\mathbf{x})\|\lesssim 1/\lambda$, which cancels out the multiplication by $\lambda$, and thus prove the following Lipschitz bound which is independent of the penalty parameter $\lambda:$
@@
$\mathcal{L}^{*}_{\lambda,i}$ is $\mathcal{O}(\ell\kappa)$-Lipschitz.
Finally, having constructed a private inexact stochastic oracle response for the smooth approximation $\mathcal{L}^{*}_{\lambda}$, we analyze an outer loop (Line 5 of Algorithm 1), showing that is optimally robust to inexact and noisy gradients. Moreover, we employ an output rule which makes use of the already-privatized iterates, thus overcoming the need of additional noise in choosing the smallest gradient. This is a subtle technical challenge which is unique to the private setting, since non-private algorithms can simply validate the gradient norm and return the minimal one, yet in our case, using the Laplace mechanism to do so would spoil the convergence rate. In particular, we show that the corresponding trajectory and output rule lead a point with small (projected-)gradient norm, as stated below:
@@
Suppose $h:\mathbb{R}^{d}\to\mathbb{R}$ is $L$-smooth, that ~$\|\widetilde{\nabla}h(\cdot)-\nabla h(\cdot)\|\leq\beta$, and consider the following update rule with $\eta=\frac{1}{2L}:$
$\mathbf{x}_{t+1}=\arg\min_{\mathbf{u}\in\mathcal{X}}\Big{\{}\big{\langle}\widetilde{\nabla}h(\mathbf{x}_{t})+\nu_{t},\mathbf{u}\big{\rangle}+\frac{1}{2\eta}\left\|\mathbf{x}_{t}-\mathbf{u}\right\|^{2}\Big{\}}~{},~{}~{}~{}\nu_{t}\sim\mathcal{N}(0,\sigma^{2}I)~{},$
with the output rule $\mathbf{x}_{\mathrm{out}}:=\mathbf{x}_{t_{\mathrm{out}}},~{}t_{\mathrm{out}}:={\arg\min}_{t\in\{0,\dots,T-1\}}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|$. If $\alpha>0$ is such that $\alpha\geq C\max\{\beta,\sigma\sqrt{d\log(T/\gamma)}\}$ for a sufficiently large absolute constant $C>0$, then with probability at least $1-\gamma:~{}\left\|\mathcal{G}_{h,\eta}(\mathbf{x}_{\mathrm{out}})\right\|\leq\alpha$ for $T=\mathcal{O}\big{(}\frac{L(h(\mathbf{x}_{0})-\inf h)}{\alpha^{2}}\big{)}$.
Notably, if $\alpha$ is predetermined, the result above allows the bias $\beta$ to be as large as $\Omega(\alpha)$ at all points, and the algorithm still reaches an $\alpha$-stationary point. Overall, applying Proposition 4.6 to $h=\mathcal{L}_{\lambda}^{*}$, we see that the (projected-)gradient norm can be as small as $\max\{\beta,\sigma\sqrt{d_{x}}\}$, up to logarithmic terms. Accounting for the smallest possible inexactness $\beta$ and noise addition $\sigma$ that ensure the the inner and outer loops, respectively, are both sufficiently private, we conclude the proof of Theorem 4.1; the full details appear in Section 8.

### v3

@@ 2.1 Setting
We impose the following assumptions which are standard in the BO literature.
@@
    $\mathcal{X}\subseteq\mathbb{R}^{d_{x}}$ is a closed convex set;
@@
    $F(\mathbf{x}_{0})-\inf_{\mathbf{x}\in\mathcal{X}}F(\mathbf{x})\leq\Delta_{F}$ for some initial point $\mathbf{x}_{0}\in\mathcal{X}$;
@@
    $f$ is twice differentiable, and $L_{1}^{f}$-smooth;
@@
    For all $\xi\in\Xi:~f(\cdot,\cdot\,;\xi)$ is $L_{0}^{f}$-Lipschitz (hence, so is $f$);
@@
    $g$ is $L_{2}^{g}$-Hessian-smooth, and for all $\mathbf{x}\in\mathcal{X}:~g(\mathbf{x},\cdot)$ is $\mu_{g}$-strongly-convex;
@@
As mentioned, these assumptions are standard in the study of BO problems and are shared by nearly all of the previously discussed works. In particular, the strong convexity of $g(\mathbf{x},\cdot)$ ensures that $\mathbf{y}^{*}(\mathbf{x})$ is uniquely defined, which is generally required in establishing the regularity of the hyperobjective. Indeed, it is known that dropping this assumption, can, in general, lead to pathological behaviors not amenable for algorithmic guarantees (cf. Chen et al. 2024; Bolte et al. 2025 and discussions therein). For the purpose of differential privacy though, the strong convexity of $g(\mathbf{x},\cdot)$ raises a subtle issue. As the standard assumption in the DP optimization literature is that the component functions are Lipschitz, which allows privatizing gradients using sensitivity arguments, strongly-convex objectives cannot be Lipschitz over the entire Euclidean space. [^3] If $g(\mathbf{x},\cdot\,;\xi)$ were Lipschitz over $\mathbb{R}^{d_{y}}$ for all $\xi\in\Xi$, then so would $g(\mathbf{x},\cdot)$, contradicting strong convexity. Therefore, strongly-convex objectives are regularly analyzed in the DP setting under the additional assumption that the domain of interest is bounded. For bilevel problems, the domain of interest for $\mathbf{y}$ is the lower-level solution set, thus we impose the following assumption.
@@
Following Assumptions 2.1 and 2.2, we denote $\ell:=\max\{L_{0}^{f},L_{1}^{f},L_{0}^{g},L_{1}^{g},L_{2}^{g}\},~\kappa:=\ell/\mu_{g}$.
@@ 3 Warm up: Privacy can leak from lower to upper level
In this section, we provide an illuminating example of privacy breaking in first-order BO, which motivates the algorithmic approach presented later in the paper. In particular, we exemplify that even when the upper-level problem supposedly does not depend on the dataset at all, the bilevel structure can give rise to hyper-gradients that are non-private.
@@
Given any dataset $S=\{\xi_{1},\dots,\xi_{n}\}\subset\mathbb{R}^{d}$, consider the BO problem corresponding to the objectives over $\mathbb{R}^{2d}:$
@@
These are smooth non-negative objectives, and $g$ is strongly-convex with respect to $\mathbf{y}$, and so we see that Assumptions 2.1 and 2.2 hold. Further note that the upper-level objective $f$ does not depend on the data $S$ at all. Nonetheless, we will show that a single hypergradient computation can break privacy.
To see this, note that for any $\mathbf{x}\in\mathbb{R}^{d}:~\mathbf{y}^{*}(\mathbf{x})=\arg\min_{\mathbf{y}}g(\mathbf{x},\mathbf{y})=\frac{1}{n}\sum_{i=1}^{n}\xi_{i}$ since the squared loss is minimized by the average. Furthermore, $\nabla_{x}f(\mathbf{x},\mathbf{y})=\mathbf{x}+\mathbf{y}$ and $\nabla^{2}_{xy}g\equiv\mathbf{0}$, so by Eq. (1) we get that $\nabla F(\mathbf{x})=\mathbf{x}+\mathbf{y}^{*}(\mathbf{x})=\mathbf{x}+\frac{1}{n}\sum_{i=1}^{n}\xi_{i}$. In particular, computing the hypergradient at the origin simply returns the dataset’s mean $\nabla F(\mathbf{0})=\frac{1}{n}\sum_{i=1}^{n}\xi_{i}$, which is well-known to be a non-private statistic of the dataset $S$ (cf. Kamath et al., 2019).
The algorithms to follow incorporate a fix to overcome such privacy leaks: instead of ever computing $\mathbf{y}^{*}(\mathbf{x})$ (as non-private algorithms do up to exponentially small precision), we estimate it using an auxiliary private method, and use the resulting private point $\widetilde{\mathbf{y}}^{*}(\mathbf{x})$ to estimate the hypergradient at $\mathbf{x}$. [^4] To be precise, in order to also avoid Hessian computations, $\nabla F(\mathbf{x})$ will be estimated by the difference of gradients at two private points close to $\mathbf{y}^{*}(\mathbf{x})$. This leads to non-negligible bias in the hypergradient estimators compared to non-private first-order BO, that we need to account for in our analysis. We further discuss this in Section 4.1.
@@ 4 Algorithm for DP bilevel ERM
In this section, we consider the bilevel ERM problem (BO/ERM), for which we denote the hyperobjective by $F_{S}$. Our algorithm is presented in Algorithm 1. We prove the following result:
@@
Assume 2.1 and 2.2 hold, and that $\alpha\leq\ell\kappa^{3}\min\{\frac{1}{2\kappa},\frac{L_{0}^{g}}{L_{0}^{f}},\frac{L_{1}^{g}}{L_{1}^{f}},\frac{\Delta_{F}}{\ell\kappa}\}$. Then there is a parameter assignment $\lambda\asymp\ell\kappa^{3}\alpha^{-1},~\sigma^{2}\asymp\ell^{2}\kappa^{2}T\log(T/\delta)\epsilon^{-2}n^{-2},~\eta\asymp\ell^{-1}\kappa^{-3},~T\asymp\Delta_{F}\ell\kappa^{3}\alpha^{-2}$, such that running Algorithm 1 satisfies $(\epsilon,\delta)$-DP, and returns $\mathbf{x}_{\mathrm{out}}$ such that with probability at least $1-\gamma:$
@@
where $K_{1}=\mathcal{O}(\Delta_{F}^{1/4}\ell^{3/4}\kappa^{5/4}),~K_{2}=\mathcal{O}(\Delta_{F}^{1/6}\ell^{5/6}\kappa^{11/6})$.
@@
Remark 4.3.
Algorithm 1 generalizes previously studied algorithms in two special cases. First, in the non-private case (when $\epsilon=\infty$ or $\delta=1$) the algorithm can skip the private inner loop and simply solve the inner problems instead in $\widetilde{\mathcal{O}}(1)$ steps, thus reducing to the first-order non-private bilevel algorithms analyzed by (Kwon et al., 2023; Chen et al., 2024, 2025). Second, if the lower-level objective is independent of $\mathbf{x}$, the partial derivative $\nabla_{x}g\equiv\mathbf{0}$ vanishes and ${\widetilde{\mathbf{y}}}_{t}\approx\arg\min_{\mathbf{y}}g(\mathbf{x}_{t},\mathbf{y})$ remains constant independently of $t$. In this case, Algorithm 1 simply reduces to DP-GD with respect to the first variable of $f$, and indeed, the $\mathcal{O}(({\sqrt{d_{x}}}/{\epsilon n})^{1/2})$ gradient bound we obtain matches the known result of DP-GD for smooth-nonconvex “single-level” optimization (Wang et al., 2017).
@@
5:  $\mathbf{x}_{t+1}=\arg\min_{\mathbf{u}\in\mathcal{X}}\left\{\left\langle\widetilde{\mathbf{g}}_{t},\mathbf{u}\right\rangle+\frac{1}{2\eta}\left\|\mathbf{u}-\mathbf{x}_{t}\right\|^{2}\right\}$ $\triangleright$ If $\mathcal{X}=\mathbb{R}^{d_{x}}$, then $\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta\widetilde{\mathbf{g}}_{t}$
@@
[…]
@@ 4.1 Analysis overview
Lemma 4.4 (Kwon et al. 2023; Chen et al. 2024, 2025).
If $\lambda\geq 2L_{1}^{f}/\mu_{g}$, then:  (i) $\left\|\mathcal{L}^{*}_{\lambda}-F\right\|_{\infty}=\mathcal{O}(\ell\kappa/\lambda)$;  (ii) $\left\|\nabla\mathcal{L}^{*}_{\lambda}-\nabla F\right\|_{\infty}=\mathcal{O}(\ell\kappa^{3}/\lambda)$;  (iii) $\mathcal{L}_{\lambda}^{*}$ is $\mathcal{O}(\ell\kappa^{3})$-smooth (independently of $\lambda$).
@@
This strategy raises a privacy concern, as we exemplified in Section 3: Since $\mathbf{y}^{*}(\mathbf{x}),\mathbf{y}^{\lambda}(\mathbf{x})$ are required in order to compute $\nabla\mathcal{L}_{\lambda}^{*}(\mathbf{x})$, and are defined as the minimizers of $g(\mathbf{x},\cdot),\mathcal{L}_{\lambda}(\mathbf{x},\cdot)$ which are data-dependent, we cannot simply compute them under DP. In other words, even deciding where to invoke the gradient oracles, can already leak user information, hence breaking privacy before the gradients are even computed. We therefore must resort to approximating them using an auxiliary private method, for which we use DP-Loc(alized)-GD (Algorithm 2). [^5] We can replace the inner solver by any DP method that guarantees with high probability the optimal rate for strongly-convex objectives, as we further discuss in Appendix B.
In our analysis, we crucially rely on the fact that $g(\mathbf{x},\cdot)$ $\mathcal{L}_{\lambda}(\mathbf{x},\cdot)$ are both strongly-convex, which implies that optimizing them produces ${\widetilde{\mathbf{y}}}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t}$ such that the distances to the minimizers, namely $\|{\widetilde{\mathbf{y}}}_{t}-\mathbf{y}^{*}(\mathbf{x}_{t})\|,\|{\widetilde{\mathbf{y}}^{\lambda}_{t}-\mathbf{y}^{\lambda}(\mathbf{x}_{t})}\|$, are small. The distance bound is key, since then Eq. (3) allows using the smoothness of $f,g$ to translate the distance bounds into an inexact (i.e., biased) gradient oracle for $\nabla\mathcal{L}_{\lambda}^{*}(\mathbf{x}_{t})$, computed at the private points ${\widetilde{\mathbf{y}}}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t}$. We are able to bound the gradient bias induced by privatizing the inner solvers, as follows:
@@
If $\lambda\geq\max\{\frac{2L_{1}^{g}}{\mu_{g}},\frac{L_{0}^{f}}{L_{0}^{g}},\frac{L^{f}_{1}}{L^{g}_{1}}\}$, then the random variables $\widetilde{\mathbf{y}}_{t},\widetilde{\mathbf{y}}^{\lambda}_{t}$ as defined in Algorithm 1 satisfy for all $t<T$ with probability at least $1-\gamma:$
@@
Bounding the Lipschitz constant of $\mathcal{L}_{\lambda,i}^{*}$ naively by applying the chain rule would result in the bound $\mathrm{Lip}(\mathbf{y}^{*})(L_{0}^{f}+2\lambda L_{0}^{g})\lesssim\lambda\mathrm{Lip}(\mathbf{y}^{*})L_{0}^{g}$, where $\mathrm{Lip}(\mathbf{y}^{*})$ is the Lipschitz constant of $\mathbf{y}^{*}(\mathbf{x}):\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{y}}$. Unfortunately, this bound grows with the penalty parameter $\lambda$, which will eventually be set large, and in particular, will grow with the dataset size $n$. We therefore need to derive a more nuanced analysis that allows obtaining a significantly tighter Lipschitz bound. We achieve this by showing that $\|\mathbf{y}^{\lambda}(\mathbf{x})-\mathbf{y}^{*}(\mathbf{x})\|\lesssim 1/\lambda$, which cancels out the multiplication by $\lambda$, and thus prove the following tight Lipschitz bound:
@@
$\mathcal{L}^{*}_{\lambda,i}$ is $\mathcal{O}(\ell\kappa)$-Lipschitz independently of $\lambda$.
Finally, having constructed a private inexact stochastic oracle response for the smooth approximation $\mathcal{L}^{*}_{\lambda}$, we analyze an outer loop (update of $\mathbf{x}_{t}$), showing that is optimally robust to inexact and noisy gradients. Moreover, we employ an output rule which makes use of the already-privatized iterates, thus overcoming the need of additional noise in choosing the smallest gradient. This is a subtle technical challenge which is unique to the private setting, since non-private algorithms can simply validate the gradient norm and return the minimal one, yet in our case, using the Laplace mechanism to do so would spoil the convergence rate. In particular, we show that the corresponding trajectory and output rule lead a point with small (projected-)gradient norm, as stated below:
@@
Suppose $h:\mathbb{R}^{d}\to\mathbb{R}$ is $L$-smooth, that $\|\widetilde{\nabla}h(\cdot)-\nabla h(\cdot)\|\leq\beta$, and consider the following update rule with $\eta=\frac{1}{2L}:\mathbf{x}_{t+1}=\arg\min_{\mathbf{u}\in\mathcal{X}}\big\{\big\langle\widetilde{\nabla}h(\mathbf{x}_{t})+\nu_{t},\mathbf{u}\big\rangle+\frac{1}{2\eta}\left\|\mathbf{x}_{t}-\mathbf{u}\right\|^{2}\big\}~,~\nu_{t}\sim\mathcal{N}(0,\sigma^{2}I)$ with the output rule $\mathbf{x}_{\mathrm{out}}:=\mathbf{x}_{t_{\mathrm{out}}},~t_{\mathrm{out}}:={\arg\min}_{t\in\{0,\dots,T-1\}}\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|$. If $\alpha>0$ is such that $\alpha\geq C\max\{\beta,\sigma\sqrt{d\log(T/\gamma)}\}$ for a sufficiently large absolute constant $C>0$, then with probability at least $1-\gamma:~\left\|\mathcal{G}_{h,\eta}(\mathbf{x}_{\mathrm{out}})\right\|\leq\alpha$ for $T=\mathcal{O}\big(\frac{L(h(\mathbf{x}_{0})-\inf h)}{\alpha^{2}}\big)$.
Notably, for any $\alpha>0$, the result above allows the bias $\beta$ to be as large as $\Omega(\alpha)$ at all points, and the algorithm still reaches an $\alpha$-stationary point. Overall, applying Proposition 4.7 to $h=\mathcal{L}_{\lambda}^{*}$, we see that the (projected-)gradient norm can be as small as $\max\{\beta,\sigma\sqrt{d_{x}}\}$, up to logarithmic terms. Accounting for the smallest possible inexactness $\beta$ and noise addition $\sigma$ that ensure the the inner and outer loops, respectively, are both sufficiently private, we conclude the proof of Theorem 4.1; the full details appear in Section 8.